# Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures

### v1

@@
Abstract
The privacy-preserving federated learning schemes based on the setting of two honest-but-curious and non-colluding servers offer promising solutions in terms of security and efficiency. However, our investigation reveals that these schemes still suffer from privacy leakage when considering model poisoning attacks from malicious users. Specifically, we demonstrate that the privacy-preserving computation process for defending against model poisoning attacks inadvertently leaks privacy to one of the honest-but-curious servers, enabling it to access users’ gradients in plaintext. To address both privacy leakage and model poisoning attacks, we propose an enhanced privacy-preserving and Byzantine-robust federated learning (PBFL) scheme, comprising three components: (1) a two-trapdoor fully homomorphic encryption (FHE) scheme to bolster users’ privacy protection; (2) a novel secure normalization judgment method to preemptively thwart gradient poisoning; and (3) an innovative secure cosine similarity measurement method for detecting model poisoning attacks without compromising data privacy. Our scheme guarantees privacy preservation and resilience against model poisoning attacks, even in scenarios with heterogeneous, non-IID (Independently and Identically Distributed) datasets. Theoretical analyses substantiate the security and efficiency of our scheme, and extensive experiments corroborate the efficacy of our private attacks. Furthermore, the experimental results demonstrate that our scheme accelerates training speed while reducing communication overhead compared to the state-of-the-art PBFL schemes.
@@ 1 Introduction
Federated learning (FL) is a new distributed learning paradigm, which enables collaborative model training among a group of distributed users and a central aggregation server, while keeping users’ sensitive data local. In an ideal scenario free from distrustful entities, FL empowers users to retain complete control over their data, thereby facilitating the development of high-quality machine learning models. However, in practice, both servers and users may lack absolute trustworthiness, leading to various attacks [1, 2, 3, 4, 5, 6] against the FL framework. In recent years, the US government has been actively promoting research aimed at creating robust and privacy-preserving FL solutions for diverse tasks such as financial crime prevention [7] and pandemic response and forecasting [8]. Incentives, including substantial prizes totaling \$$\$800,000$, have been offered to spur innovation in this area. Over the years, extensive researches [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have underscored the importance of two critical security properties in FL:
@@
The existing solutions for providing privacy and defending against poisoning attacks in FL focus on two opposing directions: privacy-preserving FL solutions aim to achieve data indistinguishability, whereas defenses against distorted gradients often rely on the similarity between malicious and benign gradients to filter out malicious ones. Specifically, defenses that extract valuable statistical information largely depend on the distinguishability of the data. Therefore, simultaneously achieving both security goals in FL remains challenging. Recently, some efforts have been made to simultaneously achieve both objectives. For instance, researchers propose privacy-preserving and Byzantine-robust FL (PBFL) schemes. However, these schemes often suffer from inefficiency [19, 20, 21, 22, 23, 24] or still make privacy leakages [22, 23, 24, 25, 26, 27]. Based on server aggregation settings, we classify existing PBFL schemes into three categories: the serverless/decentralized model, the single-server model, and the two-server model. Table I presents a qualitative comparison of FL representative works with different server aggregation settings.
@@
FedSGD [28] $1$
PrivateFL [9, 10, 11, 12, 13] $0/1/2$
RobustFL [14, 15, 16, 17, 18] $0/1/2$
PBFL [22, 23, 24] 00
PBFL [19, 20, 21] $1$
PBFL [25, 26, 27] $2$
Our scheme $2$
@@
  • Notes. In the column of “Server Setting”, $0,1,$ and $2$ represent the server-less, the single-server, and the two-server models, respectively.
@@
In the serverless/decentralized model, multiple users collaborate in training a model without any central server. Blockchain techniques are commonly employed to resist model poisoning attacks in FL with the serverless setting (e.g., [22, 23, 24]) since it provides a distributed and transparent framework for storing and managing model updates in FL, thereby reducing the risk of model poisoning attacks. However, these schemes bring additional latency due to the consensus process and result in exponentially increasing communication costs among clients, thereby diminishing the efficiency of FL. Moreover, these schemes encounter challenges in privacy protection because of the openness and transparency of the data stored on the blockchain, and the introduction of privacy protection technology will bring higher storage costs and may be limited by low transaction throughput. FL schemes with the single-server model [19, 20, 21] involve a central server in the gradient aggregation process. These schemes typically employ privacy protection techniques based on multiple keys/parameters to ensure reliable privacy protection among users, such as local sensitivities of local differential privacy in [19] and individual keys of multi-key HE in [20, 21]. Despite offering privacy protection, these schemes are often inefficient. Furthermore, relying on a single server exposes it to potential single points of failure, leaving it vulnerable to attacks from external adversaries. These attacks could result in breaches of model privacy through data infringements, hacking, or leaks. FL schemes with the two-server model (e.g., [25, 26, 27]) distribute the trust of the central server in the single-server model across two non-colluding servers. These schemes present promising solutions by fortifying resilience against single points of failure and ensuring reliable privacy protection among users without necessitating multiple keys/parameters in their privacy-preserving techniques. Consequently, they hold the potential to enhance FL performance in terms of computational cost, communication burden, and final model accuracy. Nevertheless, current schemes still face challenges in effectively protecting data privacy during their secure defenses against model poisoning attacks.
@@
In light of these challenges, this paper identifies weaknesses in PBFL schemes with the two-server model and proposes a solution. The main contributions of this paper are threefold:
@@
(1) We analyze the security of the existing PBFL scheme based on the two-server setting and show that it violates privacy. Concretely, we present that the main secure computation procedure of the existing two-server-based scheme leaks significant information about the gradients to a semi-honest server. This vulnerability may allow the server to obtain clear gradients from all users, potentially compromising the confidentiality/privacy of the poisoning defense process.
@@
(3) Through experiments, we demonstrate the efficacy of our private attacks on the existing two-server-based scheme, the effectiveness of our defense against model poisoning attacks, and the efficiency of our scheme in terms of computational and communication costs. The experiment results also show that our scheme outperforms the state-of-the-art PBFL schemes.
The remainder of this paper is organized as follows. In Section 2, we introduce the related PBFL schemes. In Section 3, we introduce the system model, the threat model, and the design goals. In Section 4, we overview the two-server-based PBFL scheme ShieldFL and present our private attacks to ShieldFL. To defeat our attacks and achieve design goals, an enhanced scheme is proposed in Section 5. Section 6 analyzes the security and the complexity of the enhanced scheme. Finally, conclusions are given in Section 7. We give the necessary notations in Table II.
@@
[[]],delimited-[]delimited-[][\![$\cdot$]\!],[ [ ⋅ ] ] , $[\cdot]_{i}$ Ciphertext/partially decrypted ciphertext $W^{(t)}$ Model parameter, a flattened vector of all weight matrixes in a learning model
@@ 3.1 System Model
3.2 Attack Model
Two types of attackers are considered, each with distinct objectives: Type I attackers aim to learn local gradients to compromise privacy-sensitive information, while Type II attackers seek to poison the model to degrade the final model performance. In the system, KC is a fully trusted entity. $U_{i\in[1,n]}$ are $n$ curious users, who may be benign or malicious. The benign users are Type I attackers who are only curious about other users’ data, while malicious users, in addition to this, may also be manipulated by Type II attackers to launch model poisoning attacks. We denote the set of malicious users as $U^{*}$. $S_{1}$ and $S_{2}$ are two honest-but-curious, non-colluding servers who honestly perform the given protocol but are curious about users’ privacy. The servers are assumed to be vigilant against Type II attackers, as they stand to benefit from an accurate and robust FL model.
Type I attackers: Both the honest-but-curious servers and the users are regarded as potential Type I attackers. A curious server may infer data privacy from local gradients submitted by the users. The users may obtain the other uses’ privacy-sensitive information by eavesdropping on the communication channel between the servers and the users.
Type II attackers: Byzantine adversaries are this type of attacker who manipulate malicious users to deviate from the correct model training by sending poisonous local gradients.
@@ 3.2 Attack Model
3.3 Design Goals
4 Security Vulnerabilities of Existing PBFL Schemes
In this section, we briefly review the existing two-server-based PBFL scheme ShieldFL [26] and identify its security vulnerabilities. We begin by discussing the technical background of ShieldFL, followed by outlining its detailed processes. Finally, we present the privacy breaches observed in ShieldFL.
4.1 Technical Background
ShieldFL utilizes a two-trapdoor partial homomorphic encryption (PHE) scheme based on Paillier’s cryptosystem to protect privacy, which contains five algorithms:
  • $\mathbf{PHE.KeyGen}(\varepsilon)\rightarrow(pk,sk)$: On input of the security parameter $\varepsilon$, the key generation algorithm first generates two odd primes $p_{1}$ and $p_{2}$ such that $|p_{1}|=|p_{2}|=\varepsilon$ and sets $N=p_{1}\cdot p_{2}$. Then, it outputs the public key $pk=(N,1+N)$ and the secret key $sk=\text{lcm}(p_{1}-1,p_{2}-1)$.
  • $\mathbf{PHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: On input of a plaintext $m\in\mathbb{Z}_{N}$, the encryption algorithm encrypts $m$ with $pk$ and outputs the ciphertext $[\![m]\!]=(1+N)^{m}\cdot r^{N}\bmod N^{2}$, where $r\in\mathbb{Z}^{*}_{N}$.
  • $\mathbf{PHE.KeySplit}(sk)\rightarrow(sk_{1},sk_{2})$: Given the secret key $sk$, two secret key shares $sk_{1}$ and $sk_{2}$ are generated through randomly dividing $sk$ meeting $\sum_{i=1}^{2}sk_{i}\equiv 0\bmod sk$ and $\sum_{i=1}^{2}sk_{i}\equiv 1\bmod N.$
  • $\mathbf{PHE.PartDec}_{sk_{i}}([\![m]\!])\rightarrow[m]_{i}$: Given a ciphertext $[\![m]\!]$, a decryption share $[m]_{i}=[\![m]\!]^{sk_{i}}$ is output through partially decrypting $[\![m]\!]$ with the secret key share $sk_{i}$.
  • $\mathbf{PHE.FullDec}([m]_{1},[m]_{2})\rightarrow m$: Given the decryption share pair $([m]_{1},[m]_{2})$, the fully decrypted plaintext $m=\frac{(\prod_{i=1}^{2}[m]_{i}\bmod N^{2})-1}{N}\bmod N$ is output.
More details of the five algorithms can be referred to [26]. The two-trapdoor PHE scheme meets the additive homomorphism such that $[\![m_{1}]\!]+[\![m_{2}]\!]=[\![m_{1}+m_{2}]\!]$ and $[\![m]\!]^{k}=[\![km]\!]$.
4.2 ShieldFL Framework
ShieldFL consists of four processes:
  1. 1.
    Setup: In this process, KC and $S_{1}$ distribute keys and an initial model, respectively.
    • KC generates a public/secret key pair $(pk,sk)$ and calls $\mathbf{PHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of independent secret key shares $(sk_{1},sk_{2})$, $(sk_{s_{1}},sk_{u_{1}}),\cdots,(sk_{s_{n}},sk_{u_{n}})$. Then, KC broadcasts $pk$; distributes $sk_{1}$ and $sk_{2}$ to $S_{1}$ and $S_{2}$, respectively; and distributes $\{sk_{s_{i}}\}_{i\in[1,n]}$ and $sk_{u_{i}}$ to $S_{1}$ and $U_{i}$, respectively.
    • $S_{1}$ releases the encrypted initial model $[\![W^{(0)}]\!]$ and its decryption share $[W^{(0)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![W^{(0)}]\!])$ to $U_{i}$ for $i\in[1,n]$.
    • Each user $U_{i}$ calculates the decryption share $[W^{(0)}]_{u_{i}}$leftarrow $\mathbf{PHE.PartDec}_{sk_{u_{i}}}([\![W^{(0)}]\!])$ and then obtains the fully decrypted initial model $W^{(0)}$leftarrow $\mathbf{PHE.FullDec}$ $([W^{(0)}]_{s_{i}},[W^{(0)}]_{u_{i}})$ to update its local model.
  2. 2.
    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$. Specifically, in the $t$th iteration, for a begin user $U_{i}$, it trains the local model $W^{(t)}$ and normalizes its local gradient $g_{i}^{(t)}$, then it sends the encrypted normalized gradient $[\![g_{i}^{(t)}]\!]$ to $S_{1}$. For a malicious user $U_{i}\in U^{*}$, it launches model poisoning and sends encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]$ to $S_{1}$.
  3. 3.
    Privacy-Preserving Defense: This process mainly consists of the following three implementations.
    • Normalization judgment. $S_{1}$ determines whether it received encrypted local gradients $[\![g_{i}^{(t)}]\!]$ are normalized by implementing a normalization judgment procedure $\mathbf{SecJudge}([\![g_{i}^{(t)}]\!])$.
    • Poisonous gradient baseline finding. $S_{1}$ finds the gradient $[\![g_{*}^{(t)}]\!]$ (among the users’ local gradients) with the lowest cosine similarity as the baseline of poisonous gradients through implementing a secure cosine similarity measurement $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$, where $[\![g_{\Sigma}^{(t-1)}]\!]$ is the aggregated gradient of the last iteration.
    • Byzantine-tolerance aggregation. First, $S_{1}$ measures the cosine similarity between the poisonous baseline and each user’s local gradient as $\cos_{i*}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$ Then, $S_{1}$ sets the confidence $co_{i}\leftarrow 1\times deg-\cos_{i*}$ for $g_{i}^{(t)}$ and calculates $co\leftarrow\sum_{i=1}^{n}co_{i},co_{i}=\lfloor\dfrac{co_{i}}{co}\cdot deg\rceil$. Finally, $S_{1}$ weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
  4. 4.
    Model Update. $S_{1}$ sends $[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{PHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$. Then $U_{i}$ decrypts to obtain the aggregated gradient plaintext $g_{\Sigma}^{(t)}$ and updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate of the $t$th training iteration.
In ShieldFL, privacy leakages primarily occur during the second and third implementations, namely, the poisonous gradient baseline finding and Byzantine-tolerance aggregation, of the privacy-preserving defense process. Therefore, we next provide a detailed procedure of the secure computation involved in these two implementations, as illustrated in Fig. 2: the implementation procedure of the secure cosine similarity measurement method $\mathbf{SecCos}$. Further information on secure computations such as $\mathbf{SecJudge}$, please refer to [26]. For further information on other secure computation procedures such as $\mathbf{SecJudge}$, please refer to [26].

### v2

@@
Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures
Jiahui Wu, Fucai Luo, Tiecheng Sun, , Haiyan Wang, Weizhe Zhang This work is supported by the Key Program of the Joint Fund of the National Natural Science Foundation of China (Grant No. U22A2036) and the Major Key Project of Peng Cheng Laboratory (Grant No. PCL2023A06).J. Wu, T. Sun and H. Wang are with the New Network Department, Peng Cheng Laboratory, Shenzhen 518000, China; F. Luo is with the School of Computer Science and Technology, Zhejiang Gongshang University, Hangzhou 310018, China; W. Zhang is with the School of Cyberspace Science Faculty of Computing, Harbin Institute of Technology, Shenzhen 518055, China and with the New Network Department, Peng Cheng Laboratory, Shenzhen 518000, China. E-mail: wujh01@pcl.ac.cn; lfucai@126.com; tiechengsun@126.com; wanghy01@pcl.ac.cn; wzzhang@hit.edu.cn. Corresponding author: W. Zhang.
@@ 1 Introduction
Federated learning (FL) is a new distributed learning paradigm, which enables collaborative model training among a group of distributed users and a central aggregation server, while keeping users’ sensitive data local. In an ideal scenario free from distrustful entities, FL empowers users to retain complete control over their data, thereby facilitating the development of high-quality machine learning models. However, in practice, both servers and users may lack absolute trustworthiness, leading to various attacks [1, 2, 3, 4, 5, 6] against the FL framework. In recent years, the US government has been actively promoting research aimed at creating robust and privacy-preserving FL solutions for diverse tasks such as financial crime prevention [7] and pandemic response and forecasting [8]. Incentives, including substantial prizes totaling \$800,000\\$800,000\$ 800 , 000, have been offered to spur innovation in this area. Over the years, extensive researches [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have underscored the importance of two critical security properties in FL:
@@
The existing solutions for providing privacy and defending against poisoning attacks in FL focus on two opposing directions: privacy-preserving FL solutions aim to achieve data indistinguishability, whereas defenses against distorted gradients often rely on the similarity between malicious and benign gradients to filter out malicious ones. Specifically, defenses that extract valuable statistical information largely depend on the distinguishability of the data. Therefore, simultaneously achieving both security goals in FL remains challenging. Recently, some efforts have been made to simultaneously achieve both objectives. For instance, researchers propose privacy-preserving and Byzantine-robust FL (PBFL) schemes. However, these schemes often suffer from inefficiency [19, 20, 21, 22, 23, 24], still make privacy leakages [22, 23, 24, 25, 26, 27], or not secure for secure multiparty computation (MPC) scenarios like PPFL applications [28, 29]. Based on server aggregation settings, we classify existing PBFL schemes into three categories: the serverless/decentralized model, the single-server model, and the two-server model. Table I presents a qualitative comparison of FL representative works with different server aggregation settings.
@@
FedSGD [30] 111
PrivateFL [9, 10, 11, 12, 13] 0/1/20/1/20 / 1 / 2
RobustFL [14, 15, 16, 17, 18] 0/1/20/1/20 / 1 / 2
PBFL [22, 23, 24] 0
PBFL [19, 20, 21] 111
PBFL [25, 26, 27, 28, 29] 222
Our scheme 222
@@
  • Notes. In the column of “Server Setting”, 0,1,0,1,0 , 1 , and 222 represent the server-less, the single-server, and the two-server models, respectively.
@@
In the serverless/decentralized model, multiple users collaborate in training a model without any central server. Blockchain techniques are commonly employed to resist model poisoning attacks in FL with the serverless setting (e.g., [22, 23, 24]) since it provides a distributed and transparent framework for storing and managing model updates in FL, thereby reducing the risk of model poisoning attacks. However, these schemes bring additional latency due to the consensus process and result in exponentially increasing communication costs among clients, thereby diminishing the efficiency of FL. Moreover, these schemes encounter challenges in privacy protection because of the openness and transparency of the data stored on the blockchain, and the introduction of privacy protection technology will bring higher storage costs and may be limited by low transaction throughput. FL schemes with the single-server model [19, 20, 21] involve a central server in the gradient aggregation process. These schemes typically employ privacy protection techniques based on multiple keys/parameters to ensure reliable privacy protection among users, such as local sensitivities of local differential privacy in [19] and individual keys of multi-key HE in [20, 21]. Despite offering privacy protection, these schemes are often inefficient. Furthermore, relying on a single server exposes it to potential single points of failure, leaving it vulnerable to attacks from external adversaries. These attacks could result in breaches of model privacy through data infringements, hacking, or leaks.
FL schemes with the two-server model (e.g., [25, 26, 27, 28, 29]) distribute the trust assumed in the traditional single-server model across two non-colluding servers. These schemes are promising as they eliminate single points of failure and enhance robustness. However, even state-of-the-art schemes such as ShieldFL contain a critical privacy vulnerability that can completely undermine their guarantees: the robust aggregation process leaks the relative relationships between clients’ gradients. A single compromised client can leverage this leakage to reconstruct other clients’ gradients in clear. In ShieldFL [26], each client encrypts its gradient under a unique key, and two servers jointly perform decryption and aggregation. To ensure no server directly learns the gradients, ShieldFL masks each gradient with random noise before cosine similarity is computed for robust aggregation. However, the same noise is applied to all gradients and to the baseline gradient. This design preserves correctness after noise removal, but it also inadvertently reveals the relative offsets between gradients. If one client is malicious, it can subtract its own gradient and the shared noise from the noisy similarity results to infer the exact gradients of all other clients, thereby breaching the privacy of the entire system. This example highlights a fundamental tension in existing two-server FL schemes: the need to preserve computational correctness (e.g., accurate robust aggregation) often conflicts with data confidentiality. Recent works such as [28, 29] employ single-key homomorphic encryption (HE) to protect the confidentiality of data from multiple clients. While this approach is effective in some cases, it fails to guarantee confidentiality between clients because all ciphertexts are encrypted under the same key. Conversely, schemes based on multi-key HE can protect inter-client confidentiality but face the challenge of performing correct computations across multiple keys. For example, existing multi-key HE-based schemes like [25, 26, 27], while offering strong privacy protection, may compromise it in order to ensure computational correctness during model aggregation and poisoning defense. In ShieldFL, although encryption and noise masking are employed to hide gradients, the robust aggregation mechanism still leaks information as explained above. This vulnerability means that once a client is compromised, the system¡¯s privacy guarantees collapse.
Therefore, despite advances in secure computation, existing two-server FL schemes still struggle to simultaneously achieve strong data confidentiality and computational correctness in the presence of model poisoning attacks.
@@
In light of these challenges, this paper identifies weaknesses in PBFL schemes with the two-server model and proposes a solution. The main contributions are threefold:
@@
(1) We analyze the security of the existing multi-key HE-based PBFL scheme based on the two-server setting and show that it violates privacy. Concretely, we present that the main secure computation procedure of the existing two-server-based scheme leaks significant information about the gradients to a semi-honest server. This vulnerability may allow the server to obtain clear gradients from all users, potentially compromising the confidentiality/privacy of the poisoning defense process.
@@
(3) Through experiments, we demonstrate the efficacy of our privacy attacks on the existing two-server-based scheme, the effectiveness of our defense against model poisoning attacks, and the efficiency of our scheme in terms of computational and communication costs. The experiment results also show that our scheme outperforms the state-of-the-art PBFL schemes.
The remainder of this paper is organized as follows. In Section 2, we introduce the related PBFL schemes. In Section 3, we introduce the system model, the threat model, and the design goals. In Section 4, we overview the two-server-based PBFL scheme ShieldFL and present our privacy attacks to ShieldFL. To defeat our attacks and achieve design goals, an enhanced scheme is proposed in Section 5. Section 6 analyzes the security and the complexity of the enhanced scheme. Finally, conclusions are given in Section 7. We give the necessary notations in Table II.
@@
[[]],[\![$\cdot$]\!],[ [ ⋅ ] ] , $[\cdot]_{i}$ Ciphertext/partially decrypted ciphertext $W^{(t)}$ Model parameter, a flattened vector of all weight matrixes in a learning model
@@ 3.1 System Model
3.2 Threat Model
We consider a federated learning system comprising a set of $n$ clients $\{U_{i}\}_{i\in[1,n]}$ and two non-colluding servers $S_{1},S_{2}$. The servers collaboratively manage encrypted aggregation and filtering tasks, while the clients periodically upload their local updates. We assume the presence of malicious clients and at most one corrupted server, a widely accepted assumption in dual-server frameworks [31, 39]. To capture different adversarial behaviors, we distinguish between honest-but-curious and malicious corruption types:
  • Honest-but-curious (semi-honest) parties follow the protocol specification but attempt to infer private information from the data they receive.
  • Malicious parties may arbitrarily deviate from the protocol, drop or fabricate messages, or tamper with the computation.
We analyze the following two adversarial scenarios:
Case 1: Semi-honest server with malicious clients. In this setting, both servers behave semi-honest, while some clients may be malicious. The semi-honest servers does not deviate from the protocol but attempts to learn sensitive information. Under this threat model, our scheme ensures:
  • The privacy of honest clients’ local updates (e.g., gradients or model parameters) is preserved.
  • The correctness of server-side computations, including aggregation and filtering, is guaranteed.
  • Malicious clients cannot compromise the privacy of others or inject invalid influence beyond their own inputs.
Case 2: Malicious server with malicious clients. In this stronger setting, one of the servers may behave maliciously, while the other remains honest-but-curious. The corrupted server may alter, drop, or fabricate computation results. Nevertheless, as long as the two servers do not collude, our scheme provides:
  • Privacy protection of honest clients’ updates remains intact due to cryptographic separation and dual control.
  • However, the correctness of computation cannot be unconditionally guaranteed, as the malicious server may tamper with intermediate results or deviate from the protocol.
This trade-off is common in dual-server designs [31, 39]. Similar to SAFEFL [39], our framework is compatible with integrating verifiable computation techniques (e.g., ZKP, MAC-based audits) to ensure correctness even in the presence of a malicious server. As verifiability introduces additional complexity and is orthogonal to our core design, we leave its implementation as future work.
3.3 Design Goals
4 Overview of Prior Work
In this section, we briefly review the existing two-server-based PBFL scheme, ShieldFL [26], and identify its security vulnerabilities. ShieldFL employs a two-trapdoor partial homomorphic encryption (PHE) scheme based on Paillier’s cryptosystem to preserve data privacy. The specific construction of the two-trapdoor PHE is provided in Appendix A. ShieldFL consists of the following four processes:
[…]
2) Local Training: Each benign user $U_{i}$ trains its local model $W^{(t)}$, encrypts its local gradient $g_{i}^{(t)}$, and sends the encrypted gradient $[\![g_{i}^{(t)}]\!]\leftarrow\mathbf{PHE.Enc}_{pk}(g_{i}^{(t)})$ to $S_{1}$. A malicious user $U_{i}\in U^{*}$ may launch model poisoning and send encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]$ to $S_{1}$.
3) Privacy-Preserving Defense: This process mainly consists of three implementations. a) Normalization judgment. $S_{1}$ determines whether it received encrypted local gradients $[\![g_{i}^{(t)}]\!]$ are normalized. b) Poisonous Gradient Baseline Finding. $S_{1}$ finds the gradient $[\![g_{*}^{(t)}]\!]$ (among the users’ local gradients) with the lowest cosine similarity as the baseline of poisonous gradients through implementing a secure cosine similarity measurement $\cos_{i\Sigma}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!])$ in Fig. 2, where $[\![g_{\Sigma}^{(t-1)}]\!]$ is the aggregated gradient of the last iteration. c) Byzantine-Tolerance Aggregation. First, $S_{1}$ measures the cosine similarity between the poisonous baseline and each user’s local gradient as $\cos_{i*}\leftarrow\mathbf{SecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$ Then, $S_{1}$ sets the confidence $co_{i}\leftarrow 1\times deg-\cos_{i*}$ for $g_{i}^{(t)}$ and calculates $co\leftarrow\sum_{i=1}^{n}co_{i},co_{i}=\lfloor\dfrac{co_{i}}{co}\cdot deg\rceil$. Finally, $S_{1}$ generates a global gradient ciphertext $[\![g_{\Sigma}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
4) Model Update. $S_{1}$ and $U_{i}$ interact to perform $\mathbf{PHE.PartDec}$ and $\mathbf{PHE.FullDec}$, so as $U_{i}$ obtains the global gradient $g_{\Sigma}^{(t)}$ and updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate of the $t$th training iteration.